Please help!
Evaluate the expression 12! using a calculator.
---
8!
Please explain how to do this!

Found 2 solutions by swincher4391, solver91311:
Answer by swincher4391(1107)

If your calculator has a ! button, it's usually under the probability section.
If you just have a four function calculator then n! = n*(n-1)*(n-2)...*1
So, we can evaluate 12! as:
12*11*10*9*8*7*6*5*4*3*2*1
and 8! = 8*7*6*5*4*3*2*1
Then you see that we can cancel out 8*7*6*5*4*3*2*1 to get
12*11*10*9 = 11880

Answer by solver91311(24713)



The first thing to do is to simplify the expression.

Note that , so . You should be able to take it from there on your calculator.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism